Problem 2 (20 points) For the given circuit, calculate Rt and It. a 1322 NA 24 31022 2022 100 V A 30 { 5022
Answer #1

Key : 132 ^^ cum 2013 3 102 202 NA - - R2 3502 & Bor / - NR3 600 20X30 = 612 100 20+30+50 Ry = 20x50 2 - 100 20+30+50 R₂ = 15132 201) zor parallel. A R3=1502 201130 - 20x30 20 + 30 = 20X30 132 ana 50 = 122 13 12r Rot us om 152 - R9 = 13+ 12+15 = 402Star to Detta and Della to oorn ea Ra Rc 7 RatRb the 0 fRz 36 2= RaRb RatRb tre R₂ = RRC R 1 R2 = R6 Rc If they are c t = had